# Rethinking Detection Based Table Structure Recognition for Visually Rich Documents

### 01 Dec 2023 v1

@@ Abstract
Table Structure Recognition (TSR) is a widely discussed task aiming at transforming unstructured table images into structured formats, such as HTML sequences. One type of popular solution is using detection models to detect components of a table, such as columns and rows, then applying a rule-based post-processing method to convert detection results into HTML sequences. However, existing detection-based studies often have the following limitations. First, these studies usually pay more attention to improving the detection performance, which does not necessarily lead to better performance regarding cell-level metrics, such as TEDS, after the post-processing. Second, some solutions over-simplify the problem and can miss some critical information. Lastly, even though some studies defined the problem to detect six components, including table, row, column, spanning cell, projected row header, and column header, to provide as much information as other types of solutions, these studies ignore the fact this problem definition is a multi-label detection because row, projected row header and column header can share identical bounding boxes. Besides these limitations, there is often a performance gap between two-stage and transformer-based detection models regarding the structure-only TEDS, even though they have similar performance regarding the COCO metrics. Therefore, in this study, we revisit the limitations of existing detection-based solutions, compare two-stage and transformer-based detection models, and identify the key design aspects for the success of a two-stage detection model for the TSR task, including the multi-class problem definition, the aspect ratio for anchor box generation, and the feature generation of the backbone network. Without proposing sophisticated new components, we applied simple methods to improve these aspects of the Cascade R-CNN model, achieved state-of-the-art performance, and improved the baseline Cascade R-CNN model by 19.32%, 11.56% and 14.77% regarding the structure-only TEDS on SciTSR, FinTabNet, and PubTables1M datasets.
@@ 1 Introduction
Figure 1: Generic workflow of transforming unstructured tabular data from document images into a structured format. Figure (a) is a sample document containing a table and other text areas. (b) represents the table detection step, and the output of this step is the table area. Note that the samples in (a) and (b) are only parts of the original results due to the page size limitation. (c) represents the step of table structure recognition. Following this study’s definition, its outputs are the detected Tables, Columns, Rows, Spanning Cells, and Headers. (d) represents the post-processing step reconstructing the complex table structure based on the detected table components. (e) is the OCR tool to extract text based on the reconstructed structure to fully transform the table image into a structured format. Notably, step (d) can also generate the cell boxes as the inputs of step (e).
(a) Four types of defined table components in the study in fernandes2023tablestrrec . This definition can infer spanning cells but cannot provide header information. Besides, the projected row headers are treated as rows, which can lead to a wrong structure.
(b) Four types of defined table components in the study in xiao2022efficient . This definition cannot provide header information, and the projected row headers are treated as rows, which can lead to a wrong structure.
(c) Six types of defined table components in the study in smock2022pubtables . The defined projected row headers in this sample share identical bounding boxes with two corresponding rows.
Figure 3: Four types of defined table components. In this sample, the defined column header shares an identical bounding box with a defined row.
@@
Datasheets are widely used in business scenarios and on the internet because of their user-friendliness to human readers. However, these documents are usually not structured, which means that it is hard to extract useful information from them and further analyze the semantic meaning automatically. Besides the unstructured format of the document, tables in these documents, which are widely used to summarize critical information, often have very complex structures and layouts, making it challenging to interpret and analyze these tables. Therefore, there have been many studies of Table Detection (TD), Table Structure Recognition (TSR), Text Detection, and Optical Character Recognition (OCR) as steps of interpreting and analyzing tables from visually rich documents. Figure 1 shows a typical workflow of extracting tables from documents and transforming the extracted tables into an HTML format, consisting of a TD, a TSR, and an OCR steps. Considering that this study formulates the TSR as an Object Detection problem, a post-processing step is required after the TSR detection model, as shown in Figure 1(d). Since there have been many studies discussing Table Detection xiao2023revisiting ; xiao2023table ; fernandes2021tabledet and OCR tools mmocr2021 ; xiao2023multi , achieving promising results, we focus on the TSR in this study, which aims at transforming the tables from visually rich documents into structured formats.
@@
TSR studies can roughly be categorized into three groups based on their problem formulations: detection-based models, image-to-sequence models, and graph-based models. Image-to-sequence models usually follow the encoder-decoder architecture and directly generate structured outputs, such as HTML sequences. Some image-to-sequence models visapp23namly ; ye2021pingan also integrate the OCR task into the model to make the model end-to-end without using extra OCR tools to extract text contents from the images. However, since these models use auto-regressive decoders, they often suffer from error accumulation shen2023divide , and their OCR capacity usually cannot generalize well because of the limitation of training data. On the other hand, graph-based models usually use segmentation or detection methods to extract table cells and treat extract table cells as nodes of a graph and further build the relation among the graph nodes. This graph-based definition makes it easier to deal with the scenarios in which table images are collected from the wild, such as rotated, distorted tables. However, graph-based models introduce extra complexity because of building graph models compared with detection-based models. By contrast, detection-based models provide an alternative way that is more straightforward to detect the table components directly and post-process the detection results with a deterministic rule-based method for reconstructing the table structure. However, the detection-based method can fail to deal with rotated and distorted samples. To sum up, these different types of approaches have their benefits and must be selected based on the application scenarios. This study focuses on the tables from well-formatted documents and proposes a detection-based solution.
@@
There have been some studies siddiqui2019deeptabstr ; hashmi2021guided ; fernandes2023tablestrrec ; xiao2022efficient ; smock2022pubtables using detection models together with a post-processing method to solve the TSR task. However, existing studies either over-simplify the problem or define a multi-label detection problem, which is challenging for two-stage object detectors. For example, studies siddiqui2019deeptabstr ; hashmi2021guided formulate the problem as only detecting columns and rows, making them hard to deal with spanning cells. TableStrRec fernandes2023tablestrrec further extend the formulation by defining regular column, regular row, irregular column, and irregular row so that the spanning cell can be inferred from these four types of components, as shown in Figure 1(a). study in xiao2022efficient simplifies the formulation of study in smock2022pubtables , directly detecting table, column, row, and spanning cell, as shown in Figure 1(b). These three formulations treat the projected row header as a regular row and do not define headers, resulting in over-simplified table structures. By contrast, study in smock2022pubtables defines six types of components, including table, column, row, spanning cell, column header, and project row header, as shown in Figure 1(c), which can provide as much structure information as other types of TSR models. However, study in smock2022pubtables does not consider that some column headers and projected row headers can share identical bounding boxes with corresponding rows. For example, as shown in Figure 1(c), the bounding boxes of the project row header can also be classified as rows. Similarly, as shown in Figure 3, a column header’s bounding box is also the row’s bounding box. Therefore, the problem definition of study in smock2022pubtables is a multi-label detection problem. This multi-label definition is challenging for typical two-stage models, such as Faster R-CNN ren2015faster and Cascade R-CNN Cai2018CascadeRD , because the classifier for the classification task is for single-label classification problems, and the feature fed into the classifier are extracted from the backbone network instead of learned queries in Sparse R-CNN sun2021sparse and DETR carion2020end . Therefore, this study follows the definition of study in smock2022pubtables , analyzes the impact of multi-class definition to the Cascade R-CNN and Sparse R-CNN, which are typical two-stage and transformer-based detectors, and further refactors this multi-label definition by extending the overlapped samples to pseudo-classes to improve the performance of our proposed Cascade R-CNN based solution. Besides, the defined table components can have extremely large aspect ratios, and the ground truth bounding boxes contain a large number of white areas that cannot generate discriminate features for the classification and regression task, as shown in Figures 1(c) and  3. Therefore, suitable aspect ratios for the anchor generation of two-stage detectors and discriminate, long-range features from the backbone network are two critical aspects for the success of a two-stage detector for the TSR. Therefore, this study tunes the aspect ratio in the Region Proposal Network (RPN), applies deformable convolution, and introduces a Spatial Attention Module that can be inserted into the backbone network to enrich long-range and context features. To sum up, the contributions of this study are four-fold:
@@
    This study revisits popular detection-based TSR definitions and formulates the TSR problem as detecting six components: table, column, row, spanning cell, projected row header, and column header, following the definition of the study in smock2022pubtables . This formulation is a multi-label detection problem, which is challenging for two-stage detectors. Therefore, this study extends the formulation to conventional single-label detection problems by generating pseudo-classes.
@@
    This study identifies and analyzes critical design aspects for the success of a TSR model, including the aspect ratios in the anchor box generation and the discriminate feature generation, then further proposes TSRDet, a Cascade R-CNN based solution, by adjusting the ratio aspects applying the deformable convolution and introducing a spatial attention module to build the long-range dependencies and context information in the backbone network.
@@
    We conduct comprehensive experiments to evaluate our proposed solution on various datasets, including SciTSR chi2019complicated , FinTabNet zheng2020global , PubTabNet zhong2020image and PubTables1M smock2022pubtables with both detection metrics and cell-level metrics. The experimental results show that our proposed TSRDet can outperform state-of-the-art models in terms of both cell-level evaluation metrics and detection metrics.
@@
    We compare the two-stage and transformer-based detectors for the TSR problem, using the TSRDet and Sparse R-CNN as two examples, and further analyze the impact of the identified critical design aspects on the transformer-based detectors.
The rest of this paper is organized as follows: Section 2 discusses related studies, including studies in Object Detection Models and Table Structure Recognition models. Section 3 describes our proposed solution. Section 4 shows the experimental results and discusses the design aspects of the proposed method. At last, we draw our conclusion and possible directions in Section 5.
@@ 2.1 Object Detection Models
On the other hand, some studies sun2021sparse ; zhang2023dense categorize the popular detectors from the perspective of Non-maximum Suppression (NMS), which is widely used to reduce the redundant predictions from the detectors. From this perspective, popular detectors can be categorized into end-to-end and none end-to-end models based on whether NMS is needed. DETR carion2020end is a typical end-to-end detector introducing transformer architecture vaswani2017attention , set prediction loss, and one-to-one label assignment to the object detection problem. Sparse R-CNN sun2021sparse further refactors the DETR model and proposes to use sparse learnable regional proposals to replace dense regional proposals and utilize a dynamic instance interactive head to regress and classify the proposals in an iterative manner. study in sun2021makes analyses the success of end-to-end detectors and argues that the one-to-one label assignment method in end-to-end detectors contributes to the success of the end-to-end models but is not sufficient to fully remove the NMS from the pipeline. This study further points out that the classification cost in the matching cost when applying one-to-one label assignment plays a key role in the success of these end-to-end models. study in zhang2023dense further analyzes combinations of the label assignment methods and queries and argues that sparse queries with one-to-one label assignment can degrade the recall, and dense queries with one-to-one label assignment are hard to optimize. To address these issues, study in zhang2023dense proposes a dense distinct queries (DDQ) method to select distinct queries from dense queries using a class-agnostic NMS, achieving promising precision and recall. It is worth mentioning that these end-to-end detector can easily be extended to none end-to-end solutions by adapting many-to-one label assignments hong2022dynamic and NMS.
@@ 2.2 Table Structure Recognition
There have been many studies Chi2019ComplicatedTS ; Adiga2019TableSR ; zheng2021global ; xiao2022efficient ; Schreiber2017DeepDeSRTDL ; liu2022neural discussing the TSR problem in recent years. As mentioned earlier in Section 1, we can roughly categorize these solutions into image-to-sequence, detection-based, and graph-based models. Image-to-sequence based models usually define the ground truth as structured sequences, such as HTML sequences, built on the transformer architecture vaswani2017attention , and follow an encoder-decoder architecture. For instance, TableMaster ye2021pingan is a typical image-to-sequence based model that can generate HTML sequences. More specifically, TableMaster follows the architecture of MASTER lu2021master , which is originally designed for the scene text generation following the transformer architecture vaswani2017attention , and further improved the encoder part by introducing a Multi-Aspect Global Context Attention. In contrast, TableMaster has two branches designed for the HTML sequence generation and bounding box regression. Similarly, MTL-TabNet visapp23namly also follows the encoder-decoder architecture but contains three decoders for the cell box regression, cell content recognition, and HTML sequence generation. DRCC shen2023divide argues that the error accumulation problem degrades the performance of image-to-sequence TSR models, especially when the input image is large. Therefore, DRCC proposes a two-step decoder architecture, which first decodes the input image into rows and then decodes the rows in cell sequences. VAST huang2023improving pays more attention to the imprecise bounding boxes of table cells and proposes a Coordinate Sequence Decoder to improve the model’s ability to generate accurate bounding boxes and introduces a visual-alignment loss to align the visual and structural information. To sum up, this type of method is usually based on the encoder-decoder architecture and can be trained end-to-end without using post-processing methods.
@@
On the other hand, detection-based models usually define the problem as detecting different table components and applying a post-processing method to reconstruct table structures. DeepTabStR siddiqui2019deeptabstr proposes to detect columns and rows to obtain the table cells. However, DeepTabStR ignores the row/column span in the tables, which means that it cannot recover the hierarchical structures of tables. TableStrRec Fernandes2023TableStrRecFF extends the DeepTabStR, defining four types of table components: regular columns, irregular columns, regular rows, and irregular rows. Then, the spanning cells across multiple columns can be inferred from the difference between the regular and irregular columns when they are overlapped, and the spanning cells across multiple rows can be inferred from regular and irregular rows similarly. PubTables-1M smock2022pubtables is another typical detection-based approach that defines six table components: table, column, row, spanning cell, projected row header, and columns header, in which projected row header and columns header are for the function analysis and other components can be used to reconstruct the complex table structure.

### 10 Jan 2024 v2

# Rethinking Detection Based Table Structure Recognition for Visually Rich Document Images

@@ Abstract
Table Structure Recognition (TSR) is a widely discussed task aiming at transforming unstructured table images into structured formats, such as HTML sequences, to make text-only models, such as ChatGPT, that can further process these tables. One type of solution is using detection models to detect table components, such as columns and rows, then applying a rule-based post-processing method to convert detection results into HTML sequences. However, existing detection-based models usually cannot perform as well as other types of solutions regarding cell-level TSR metrics, such as TEDS, and the underlying reasons limiting the performance of these models on the TSR task are also not well-explored. Therefore, we revisit existing detection-based models comprehensively and explore the underlying reasons hindering these models’ performance, including the improper problem definition, the mismatch issue of detection and TSR metrics, the characteristics of detection models, and the impact of local and long-range features extraction. Based on our analysis and findings, we apply simple methods to tailor a typical two-stage detection model, Cascade R-CNN, for the TSR task. The experimental results show that the tailored Cascade R-CNN based model can improve the base Cascade R-CNN model by 16.35% on the FinTabNet dataset regarding the structure-only TEDS, outperforming other types of state-of-the-art methods, demonstrating that our findings can be a guideline for improving detection-based TSR models and that a purely detection-based solution is competitive with other types of solutions, such as graph-based and image-to-sequence solutions.
@@
{highlights}
Revealing the issues of detection-based Table Structure Recognition (TSR) studies.
Analysis and explanation of critical aspects of detection models and the TSR task.
Proposing a Cascade R-CNN based solution, based on the analysis.
Achieving STOA performance on various datasets regarding COCO and TEDS metrics.
Further explanation on the impact of critical aspects based on experimental results.
@@ 1 Introduction
Datasheets and scanned documents are widely used in business scenarios and on the internet because of their user-friendliness to human readers. However, these documents are usually not structured, which means that it is hard to extract useful information from them and further analyze the semantic meaning automatically. Besides the unstructured format of the document, tables in these documents, which are widely used to summarize critical information, often have very complex structures and layouts, making it challenging to interpret and analyze these tables. Therefore, there have been many studies rastan2019texus ; yildiz2005pdf2table ; mendes2017tabula ; fernandes2023tablestrrec ; xiao2022efficient ; xiao2022handling ; zhong2020image ; zheng2021global ; qiao2021lgpma ; ma2023robust ; li2022table ; huang2023improving ; nassar2022tableformer ; visapp23namly discussing the Table Structure Recognition (TSR) problem aiming at transforming unstructured table images into structured formats, such as HTML sequences, so that the transformed structured tabular data can be further analyzed by text-only models, such as ChatGPT openai_chatgpt . Even though some studies and tools rastan2019texus ; yildiz2005pdf2table ; mendes2017tabula can parse the Portable Document Format (PDF) files directly to extract the content and tables, these solutions cannot deal with scanned documents and table images and their performance is often limited on the TSR problem. Therefore, this study focuses on the deep learning based models.
@@
TSR studies can roughly be categorized into three groups based on their problem formulations: detection-based models, image-to-sequence models, and graph-based models. Image-to-sequence models usually follow the encoder-decoder architecture and directly generate structured outputs, such as HTML sequences. Some image-to-sequence models visapp23namly ; ye2021pingan also integrate the OCR task into the model to make the model end-to-end without using extra OCR tools mmocr2021 ; xiao2023multi to extract text contents from the images. However, since these models use auto-regressive decoders, they often suffer from error accumulation shen2023divide , and their OCR capacity usually cannot generalize well because of the limitation of training data. On the other hand, graph-based models usually use segmentation or detection methods to extract table cells, treat extract table cells as nodes of a graph, and further build the relation among the graph nodes. This graph-based definition makes it easier to deal with the scenarios in which table images are collected from the wild, such as rotated, distorted tables. However, graph-based models introduce extra complexity because they need to build extra graph models compared with detection-based models. By contrast, detection-based models are more straightforward in detecting the table components directly and post-processing the detection results with a deterministic rule-based method for reconstructing the table structure. However, detection-based methods can fail to deal with rotated and distorted samples. Besides, detection-based models usually cannot perform as well as other types of solutions regarding cell-level TSR metrics, such as TEDS zhong2020image . Therefore, these different types of approaches have their benefits and must be selected based on the application scenarios. In this study, we focus on the table images from well-formatted documents and explore the underlying reasons hindering the performance of detection-based models.
@@
There have been many studies siddiqui2019deeptabstr ; hashmi2021guided ; fernandes2023tablestrrec ; xiao2022efficient ; smock2022pubtables using detection models together with a post-processing method to solve the TSR task. However, existing studies either over-simplify the problem or define a multi-label detection problem, which is challenging for two-stage object detectors. For example, some studies siddiqui2019deeptabstr ; hashmi2021guided ; fernandes2023tablestrrec ; xiao2022efficient do not define Column Header as detection target, making it impossible to provide information regarding the header cells. By contrast, PubTables1M smock2022pubtables defines six types of components, including Table, Column, Row, Spanning cell, Column Header, and Projected Row Header, which can provide as much structure information as other types of TSR models. However, PubTables1M smock2022pubtables does not consider that some Column Headers and Projected Row Headers can share identical bounding boxes with corresponding Rows, making this definition a multi-label detection task. Besides these issues of problem formulation, detection models used in these studies are trained to optimize their detection performance. However, since the complex structures of tables are processed by a post-processing step inferring defined table components, such as Columns and Rows, a model with good detection performance cannot necessarily lead to good performance in TSR metrics, such as TEDS. Moreover, some critical characteristics of detection models are not considered in the model design and problem formulation in existing studies. For example, typical two-stage detection models, such as Faster R-CNN ren2015faster and Cascade R-CNN Cai2018CascadeRD , are not suitable for multi-label detection tasks, while transformer-based detection models, such as DETR carion2020end and Sparse R-CNN sun2021sparse , can achieve promising results on multi-label detection tasks. Another example is that for two-stage detection models, regional proposal generation plays a crucial role in the model’s performance, and the defined components in table images have different aspect ratios compared with common objects. At last, many studies apply deformable convolution dai2017deformable to improve the models’ performance regarding detection evaluation metrics, such as COCO metric lin2014microsoft . However, simply applying deformable convolution can degrade the model’s performance regarding the TEDS, and it is necessary to extract long-range dependencies while improving the local feature extraction. Therefore, in this study, we comprehensively revisit existing detection-based solutions and further explore the possible reasons hindering the performance of detection-based models for the TSR task. Based on our findings and analysis, we apply three simple methods to a typical two-stage detection model, Cascade R-CNN, including tuning the aspect ratios and increasing the number of region proposals in regional proposal generation, transforming the multi-label detection task into the single-label task, and introducing a Spatial Attention Module to build long-range dependencies. The experimental results show that our proposed method can achieve state-of-the-art performance with very simple methods, demonstrating that our findings can be a guideline for further improvement of detection-based solutions. To sum up, the contributions of this study are four-fold:
@@
    We comprehensively revisit existing detection-based TSR models and explore possible reasons hindering the performance of these models, including the problem formulation, the mismatch of optimization target of detection models and the TSR evaluation metrics, the characteristics of detection models, and the impact of feature extraction. Our analysis and findings can be a guideline for further improving the performance of detection-based TSR models.
@@
    Based on our analysis and findings, we apply three simple methods to improve Cascade R-CNN, including proposing a pseudo-class generation method to transform multi-label detection into a regular single-label detection problem, adjusting the ratio aspects and the number of regional proposals in the region proposal generation, applying the deformable convolution and introducing a Spatial Attention Module to build the long-range dependencies and context information in the backbone network.
@@
    We conduct extensive experiments to evaluate our proposed solution on various datasets, including SciTSR chi2019complicated , FinTabNet zheng2020global , PubTabNet zhong2020image and PubTables1M smock2022pubtables with both detection metrics and cell-level TSR metrics. The experimental results show that our proposed solution can outperform state-of-the-art models in terms of detection and cell-level TSR metrics.
@@
    We further verify our analysis and findings with experiments and discuss and explain more useful insights from the experimental results for further model design.
The rest of this paper is organized as follows: Section 2 discusses related studies, including studies in Object Detection and Table Structure Recognition. Section 3 explores and discusses the reasons that hinder the performance of detection-based TSR models. Section 4 describes our proposed solution based on our analysis and findings. Section 5 shows the experimental results and discusses the design aspects of the proposed method. At last, we draw our conclusion and possible directions in section 6.
@@ 2.1 Object Detection Models
On the other hand, some studies sun2021sparse ; zhang2023dense categorize the popular detectors from the perspective of Non-maximum Suppression (NMS), which is widely used to reduce the redundant predictions from the detectors. From this perspective, popular detectors can be categorized into end-to-end and none end-to-end models based on whether NMS is needed. DETR carion2020end is a typical end-to-end detector introducing transformer architecture vaswani2017attention , set prediction loss, and one-to-one label assignment to the object detection problem. Sparse R-CNN sun2021sparse further refactors the DETR model and proposes to use sparse learnable regional proposals to replace dense regional proposals and utilize a dynamic instance interactive head to regress and classify the proposals in an iterative manner. Study sun2021makes analyses the success of end-to-end detectors and argues that the one-to-one label assignment method in end-to-end detectors contributes to the success of the end-to-end models but is not sufficient to fully remove the NMS from the pipeline. This study further points out that the classification cost in the matching cost when applying one-to-one label assignment plays a key role in the success of these end-to-end models. Study zhang2023dense further analyzes combinations of the label assignment methods and queries and argues that sparse queries with one-to-one label assignment can degrade the recall, and dense queries with one-to-one label assignment are hard to optimize. To address these issues, study zhang2023dense proposes a dense distinct queries (DDQ) method to select distinct queries from dense queries using a class-agnostic NMS, achieving promising precision and recall. It is worth mentioning that these end-to-end detector can easily be extended to none end-to-end solutions by adapting many-to-one label assignments hong2022dynamic and NMS. In this study, we refer to models using transformer architecture, set prediction loss, and their variations as transformer-based detection models, such as DETR, Sparse R-CNN and Deformable-DETR zhu2020deformable .
@@ 2.2 Table Structure Recognition
There have been many studies Chi2019ComplicatedTS ; Adiga2019TableSR ; zheng2021global ; xiao2022efficient ; Schreiber2017DeepDeSRTDL ; liu2022neural discussing the TSR problem in recent years. As mentioned earlier in Section 1, we can roughly categorize these solutions into image-to-sequence, detection-based, and graph-based models. Image-to-sequence based models usually define the ground truth as structured sequences, such as HTML sequences, built on the transformer architecture vaswani2017attention , and follow an encoder-decoder architecture. For instance, TableMaster ye2021pingan is a typical image-to-sequence based model that can generate HTML sequences. More specifically, TableMaster follows the architecture of MASTER lu2021master , which is originally designed for the scene text generation following the transformer architecture vaswani2017attention , and further improved the encoder part by introducing a Multi-Aspect Global Context Attention. Besides, TableMaster has two branches designed for the HTML sequence generation and bounding box regression. Similarly, MTL-TabNet visapp23namly also follows the encoder-decoder architecture but contains three decoders for the cell box regression, cell content recognition, and HTML sequence generation, respectively. DRCC shen2023divide argues that the error accumulation problem degrades the performance of image-to-sequence TSR models, especially when the input image is large. Therefore, DRCC proposes a two-step decoder architecture, which first decodes the input image into rows and then decodes the rows in cell sequences. VAST huang2023improving pays more attention to the imprecise bounding boxes of table cells and proposes a Coordinate Sequence Decoder to improve the model’s ability to generate accurate bounding boxes and introduces a visual-alignment loss to align the visual and structural information. To sum up, this type of method is usually based on the encoder-decoder architecture and can be trained end-to-end without using post-processing methods. Since the ground truth sequences used in image-to-sequence models usually contain information regarding spanning cells and header cells, these models can handle complex structures with spanning cells and identify header cells.
@@
On the other hand, detection-based models usually define the problem as detecting different table components and applying a post-processing method to reconstruct table structures. DeepTabStR siddiqui2019deeptabstr proposes to detect columns and rows to obtain the table cells. However, DeepTabStR ignores the row/column-span in the tables, which means that it cannot recover the hierarchical structures of tables. TableStrRec Fernandes2023TableStrRecFF extends the DeepTabStR, defining four types of table components: regular columns, irregular columns, regular rows, and irregular rows. Then, the spanning cells across multiple columns can be inferred from the difference between the regular and irregular columns when they are overlapped, and the spanning cells across multiple rows can be inferred from regular and irregular rows similarly. PubTables1M smock2022pubtables is another typical detection-based approach that defines six table components: table, column, row, spanning cell, Projected Row Header, and Column Header, in which Projected Row Header and Column Header are for the function analysis, and other components can be used to reconstruct the complex table structure. Among these formulations, only the problem formulation of PubTables1M can provide as much information as image-to-sequence models because it can provide header cell information and reconstruct the complex table structure. Besides, these detection-based models need an extra deterministic rule-based post-processing method to infer the table structure from detected table components, meaning they are not end-to-end.